OpenStudy (anonymous):

Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain.

OpenStudy (anonymous):

n an 1 9 2 3 3 −3

OpenStudy (anonymous):

common difference=t2-t1=3-9=-6

OpenStudy (anonymous):

tn=a1+(n-1) d

OpenStudy (anonymous):

an = 9 − 3(n − 1) where n ≤ 9 an = 9 − 3(n − 1) where n ≥ 1 an = 9 − 6(n − 1) where n ≤ 9 an = 9 − 6(n − 1) where n ≥ 1

OpenStudy (anonymous):

I got D?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

yw